QUESTION 3 The IUPAC name of this compound is: Br C=CH-CECH NEC-CH CH2CH2CH3
Answer #1

to NAME I select longest chain Z=CH-CE CH int=c-ne Ter-CH-CH2 CH3-CH₂-CH $ 6 -cn-c=cH -CECH ls 4 3 2 1 CEN start the numberin